OpenStudy (anonymous):

x^4-37x^2+36 find the factors

OpenStudy (anonymous):

let x^4 = t^2 Thus, t^2 - 37t + 36 = 0

OpenStudy (anonymous):

\[(x^2-36)(x^2-1)\]

OpenStudy (anonymous):

the roots are 1 and 36

OpenStudy (anonymous):

also -1 and -6

OpenStudy (anonymous):

Therefore x = \[\pm1 , \pm\]

OpenStudy (anonymous):

Sorry I meant roots for t are 1 and 36

OpenStudy (anonymous):

Therefore \[x = \pm1 , \pm 6\]

OpenStudy (anonymous):

so (x-1)(x+1)(x-6)(x+6)?

OpenStudy (anonymous):

oui

OpenStudy (anonymous):

Yup right littleice!